OpenStudy (anonymous):

Help please f(x) = cubed root 3x^3-1 find f^-1(x) and (f^-1 of) (135)

OpenStudy (asnaseer):

the easiest way of tackling problems like this is to write them in the form y=f(x), then swap the x and y's and finally solve for y. so, in this case we are given:\[y=f(x)=\sqrt[3]{3x^3-1}\]so first step is to swap the x and y's, giving us:\[x=\sqrt[3]{3y^3-1}\]now we need to solve for y. so lets cube both sides to get:\[x^3=3y^3-1\]therefore:\[y=\sqrt[3]{\frac{1+x^3}{3}}\]this is your inverse function. so now replace y with \(f^{-1}(x)\) to get:\[f^{-1}(x)=\sqrt[3]{\frac{1+x^3}{3}}\] to work out \(f^{-1}(135)\) just substitute x=135 into the equation above.

OpenStudy (anonymous):

thankyou very much!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

can i ask one more question for a different problem : f(x)=2x+3 find (f(x))^-1

OpenStudy (asnaseer):

Its best if you post this as a new question.

OpenStudy (asnaseer):

but the principal to solve it the same as I outlined for you above.